# A Watermark-Conditioned Diffusion Model for IP Protection

### v1

@@ Abstract
The ethical need to protect AI-generated content has been a significant concern in recent years. While existing watermarking strategies have demonstrated success in detecting synthetic content (detection), there has been limited exploration in identifying the users responsible for generating these outputs from a single model (owner identification). In this paper, we focus on both practical scenarios and propose a unified watermarking framework for content copyright protection within the context of diffusion models. Specifically, we consider two parties: the model provider, who grants public access to a diffusion model via an API, and the users, who can solely query the model API and generate images in a black-box manner. Our task is to embed hidden information into the generated contents, which facilitates further detection and owner identification. To tackle this challenge, we propose a Watermark-conditioned Diffusion model called WaDiff, which manipulates the watermark as a conditioned input and incorporates fingerprinting into the generation process. All the generative outputs from our WaDiff carry user-specific information, which can be recovered by an image extractor and further facilitate forensic identification. Extensive experiments are conducted on two popular diffusion models, and we demonstrate that our method is effective and robust in both the detection and owner identification tasks. Meanwhile, our watermarking framework only exerts a negligible impact on the original generation and is more stealthy and efficient in comparison to existing watermarking strategies.
@@ 1 Introduction
Figure 1: Illustration of our proposed setting. All users access the diffusion model by querying the public API and are assigned a unique watermark. The generation process is conditioned on the watermark, and each user’s generated outputs would carry specific fingerprinting information which is further utilized to identify the owner of the generated image.
@@
To enable the traceability of diffusion-generated images, a commonly employed strategy is to embed a unique fingerprint to contents generated by an individual user and then forensically identify the owner from the watermarked image. A line of previous works [1, 25, 11, 28, 42] have extensively investigated this approach within the realm of traditional multimedia copyright protection, which is commonly referred to as post-hoc watermarking. Typically, these strategies involve embedding an imperceptible fingerprint into the generative content while leaving an identifiable trace that can be detected using a pre-designed mechanism. However, post-hoc watermarking requires additional computational costs for watermark injection and is more susceptible to circumvention. For instance, in the event of model leakage, the post-processing module can be easily detected and bypassed by attackers (see Section 5.5). Recently, the Stable Signature [10] investigated the latent diffusion paradigm and proposed a method that incorporates watermarking into each latent decoder before distribution among individual users. However, solely fingerprinting the latent decoder limits their application scenario to the latent diffusion model only and can be easily circumvented by retraining the latent decoder on a clean dataset. At the same time, since only a fixed watermark could be embedded into the latent diffusion paradigm, every latent diffusion model has to be fine-tuned before being distributed to the users, making it hard to use in large-scale real-world systems. Therefore, a naturally asked question is whether it is possible to embed fingerprints during the generation process such that the user-specific information is embedded to the generative images without the need for user-specific model retraining.
@@
In this paper, as shown in Figure 1, we introduce WaDiff, a watermark-conditioned diffusion model, which incorporates the watermark as a conditioned input and generates images with unique fingerprints tailored to individual users. Unlike previous approaches [10, 34], which solely focus on fingerprinting the latent decoder in the diffusion model, our watermarking strategy is seamlessly integrated into the diffusion generation process. This makes our approach more general and applicable to other types of diffusion models, such as DDPM [13], that do not include a latent decoder and, at the same time, eliminate the need for post-processing. Our WaDiff builds upon a pre-trained diffusion model. To accommodate the inclusion of the watermark information, we make slight modifications to the input layers with the channel expanding. Specifically, we first embed the watermark bits through linear transformation and then concatenate this projected vector with the original input to construct the watermark-conditioned input. We design a unified watermarking framework with two novel objective functions named message retrieval loss and consistency loss. The message retrieval loss ensures the effective embedding of fingerprints into the generated content, allowing for successful retrieval of the embedded watermark. The consistency loss, on the other hand, ensures that the inclusion of the watermark has a negligible effect on the overall generation quality.
@@
To this end, we evaluate our method on two popular open-source diffusion models and perform both the detection and identification tasks. We also make detailed comparisons with widely used post-hoc watermarking strategies and Tree-Ring [32], a training-free fingerprinting framework for diffusion models. Our experimental results demonstrate that our efficient watermarking strategy enables accurate detection and identification among a large-scale system with million users, and is robust across various data augmentations. Moreover, generated images after watermarking could still maintain an exceptional generation quality and are visually identical. We organize our contributions as follows:
@@
  • Different from previous work, we propose to efficiently incorporate user-specific fingerprints into the diffusion sampling process without customized fine-tuning.
@@
  • We introduce WaDiff, a watermarked conditioned diffusion model, along with a unified watermarking framework that manipulates the watermark as conditioned input. Our approach allows for the effective embedding of user-specific watermarks while having a negligible impact on the original generation quality.
@@
  • Extensive experiments demonstrate that our watermarking framework achieves precise and robust performance in detecting AI-generated content and identifying the source owner of generated images from millions of users.
@@ 2.1 Diffusion Models
Diffusion models have recently shown tremendous capability for high-quality image generation [13, 29, 27, 26]. Depending on the space in which the generation process is performed, diffusion models can be divided into pixel-space diffusion models [13, 8] and latent-space diffusion models [26]. In pixel-space diffusion models, images are directly generated starting from sampled Gaussian noise [13]. To reduce the computational complexity, latent diffusion models have been proposed to first generate latent features from noise and then decode the latent features to images by VAE [26]. With the powerful generation capability, they have demonstrated amazing results on various computer vision tasks, such as text-to-image generation [22, 27], sketch-to-image generation [21], text-guided image editing [22], etc.
@@ 2.2 Image Watermarking
Image watermarking has been a broadly explored technique for decades. Traditional strategies typically start from a host image and inject watermark information directly in the spatial domain [25, 28], or through certain domain transformations such as Discrete Cosine Transform (DCT) [1] and Discrete Wavelet Transform (DWT) [11]. With advancements in deep learning, researchers have also explored the use of deep-learning techniques to replace manually designed watermark patterns. For example, a line of works [42, 18, 30, 41, 37] leverages the capability of Deep Neural Networks (DNNs) to improve the stealthiness and robustness of the watermark. In addition to facilitating image protection, image watermarking techniques have been proved useful in various security challenges, such as model ownership verification [39, 33, 38], backdoor attack [16], dataset copyright protection [15], and forensic adversarial defense [4]. In our method, we leverage image watermarking to embed unique watermark information for individual users, thereby facilitating subsequent detection.
@@ 2.3 Fingerprinting in Diffusion Models
In light of recent advancements in generative AI, researchers have been exploring the application of watermarking techniques to safeguard the generative contents [36, 35, 14, 9]. The Stable Signature [10] first proposed a watermarking scheme for latent diffusion models [26] by fine-tuning a set of latent decoders guided by a pre-trained Hidden [42] extractor. They then distributed these customized decoders to individual users for both detection and identification. Building upon this work, [34] improved the scalability by incorporating a message-matrix into the latent decoder, allowing multiple watermarks to be carried within a single architecture. However, their methods only focus on manipulating the latent decoder while keeping the diffusion process intact, which makes their methods only applicable to latent-based diffusion models. Also, this leaves them vulnerable to attacks like simply retraining the fingerprinted latent decoder on a clean dataset. In a parallel work called Tree-Ring [32], researchers concealed the fingerprint within the frequency domain of the initial noisy vector and performed detection by reversing the generative image back to the noisy vector. They then compared it with the original noisy pattern to determine the presence of the watermark. Despite its training-free nature, the detection process heavily relies on the time-consuming reversion process with deterministic samplers like DDIM [29].
In contrast to previous studies, we propose the WaDiff along with a simple and unified watermarking framework for owner identification. Our method seamlessly integrates the fingerprinting process into the image generation process, which can be applied with various diffusion types and sampling schedules, resulting in enhanced stealthiness and compatibility.
Note that other research works have developed fingerprinting techniques on diffusion models for network watermarking [40] and data protection [20, 6]. These topics are beyond our scope, and will not be discussed in this paper.
@@ 3 Problem Setting
The model provider deploys a generation model and grants public access to $m$ registered users. Considering the intellectual property (IP) protection, the model architecture and parameters are concealed from users, which means only black-box access is permitted and all the internal information remains encrypted. Note that the black-box scenario is quite common in practice, and has been widely adopted in current generative models such as the Midjourney and ChatGPT [3]. Each user denoted as $u_{i}$, is required to register before usage and would be then assigned a unique ID $i$ where $i\in\{1,\dots,m\}$, representing the user’s unique identity. One of the users $u_{i}$ would use the provided generation model to generate a picture $\mathbf{p}$. In practice, the $\mathbf{p}$ might be further processed under several image manipulations $f$ such as resizing and compression. We aim to build a security auditing system to enable the detection and tracing of diffusion-generated content $\mathbf{p}$.
We divided our task into two challenges. The first challenge involves determining whether the generative content originates from our diffusion model, which we refer to as detection. However, in scenarios where multiple users access the model, simple detection of the generative content is insufficient to differentiate who generated a particular image. Therefore, our second challenge focuses on a more complex scenario, which entails accurate identification of the owner of $f(\mathbf{p})$, denoted as $u$, from a pool of users. To address both challenges effectively, we propose a unified watermarking framework named WaDiff, which incorporates the embedding of user-specific fingerprints within the generation process. In other words, the generative images of each user are watermarked with unique information that distinguishes them from other users and facilitates both detection and identification.
@@ 4.2 Rooting Fingerprints in Diffusion Process
Pre-training Watermark Decoders
@@
Inspired by [10], we first train an image encoder $\mathcal{W}$ and decoder $\mathcal{D}$ such that $\mathcal{D}$ could retrieve the message $w$ pre-embedded by $\mathcal{W}$. The $\mathcal{W}$ is then discarded and only the pre-trained $\mathcal{D}$ is left to serve as a reference to fine-tune the diffusion model. Specifically, the encoder $\mathcal{W}$ takes an image $\mathbf{x}$ and a $n$-bit binary message $w\in\{0,1\}^{n}$ as input and encodes $w$ as an imperceptible residual $\delta$ added to $\mathbf{x}$, where the extractor $\mathcal{D}$ aims to restore the pre-encoded $w$ from the watermarked input $\mathcal{W}(\mathbf{x})=\mathbf{x}+\delta$. To train $\mathcal{W}$ and $\mathcal{D}$ in an end-to-end manner, we utilize two loss terms to fulfill the optimization objective. First, to accurately restore $w$ from $\mathcal{W}(\mathbf{x})$, we minimize the Binary Cross Entropy (BCE) between the decoded output $\mathcal{D}(\mathcal{W}(\mathbf{x}))$ and ground-truth $w$. Second, to reduce the visibility of $\delta$, we penalize $\|\delta\|_{2}$ such that the added message perturbation is less perceptible. We further notice that the $\mathcal{D}$ obtained with the above training schedule is sensitive to several data manipulations such as resizing and compression. Therefore, similar to the simulation layer in [10], we transform $\mathcal{W}(\mathbf{x})$ with random data augmentations during the training stage to improve the robustness of the watermark system. We formulate the complete training objective as follows:
$\min_{\mathcal{W},\mathcal{D}}\mathcal{L}_{BCE}(\mathcal{D}(f(\mathcal{W}(\mathbf{x}))),w)+\gamma\|\delta\|_{2},$ (3)
@@
where $\gamma>0$ is a hyperparameter to control the visibility of $\delta$ and $f$ is an image transformation function sampled from a pool of data augmentations.
Watermark-conditioned Diffusion Model
@@
In contrast to post-hoc strategies, where the fingerprinting process occurs after the entire generation process, our approach imprints fingerprints during the sampling process. To discern the source user of generated output, we assign each user $u_{i}$ a unique $w_{i}$ as the conditioned input of the diffusion model such that $w_{i}$ could be embedded to the generated content during inference and correctly restored by the pre-trained $\mathcal{D}$.
@@
To incorporate the watermark $w_{i}$ into the generation process, we expand the input channels at each denoising step $t$. This is achieved by applying a linear layer to project $w_{i}$ into $\mathcal{P}(w_{i})\in\mathbb{R}^{\widetilde{C}\times H\times W}$ and then concatenating it with the original latent variable $\mathbf{x}_{t}\in\mathbb{R}^{C\times H\times W}$ along their first dimension. We denote the conditioned latent as $\mathbf{\hat{x}}_{t}=concat(\mathbf{x}_{t},\mathcal{P}(w_{i}))\in\mathbb{R}^{(C+\widetilde{C})\times H\times W}$. Note that channel expansion is a common manipulation to integrate additional information and has been widely used in previous works [42, 30]. The next step involves embedding $w_{i}$ into the generative content to allow for detection by the pre-trained decoder $\mathcal{D}$. However, the sampling process usually involves multiple denoising steps to obtain the generative output, which makes it hard to incorporate into the fine-tuning process. To address this issue, we start by restoring the original image $\mathbf{x}_{0}$ within a single step. Revisiting the forward noising in Equation 2, $\mathbf{x}_{0}$ could be directly recovered from $\mathbf{x}_{t}$ by subtracting the second noise term and subsequently rescaling. In the absence of the ground truth $\epsilon$, we take the prediction from the diffusion model as an estimate. In our watermark-conditioned model, we replace $\epsilon$ with $\epsilon_{\theta}(\hat{\mathbf{x}}_{t};w_{i})$ and construct the conditioned reverse of $\mathbf{x}_{0}$ as:
$\hat{\mathbf{x}}_{0}^{t}=\frac{\mathbf{x}_{t}-\sqrt{1-\overline{\alpha}_{t}}\epsilon_{\theta}(\hat{\mathbf{x}}_{t};w_{i})}{\sqrt{\overline{\alpha}_{t}}}.$ (4)
@@
$\min_{\theta}\mathcal{L}_{m}(\mathcal{D}(\hat{\mathbf{x}}_{0}^{t}),w_{i}).$ (5)
By optimizing Equation 5, we ensure that the pre-embedded $w_{i}$ in $\hat{\mathbf{x}}_{0}^{t}$ could be detected by $\mathcal{D}$. However, the image quality of $\hat{\mathbf{x}}_{0}^{t}$ is significantly influenced by the denoising step $t$, resulting in higher image quality for smaller values of $t$ and noisier images for larger values of $t$. Therefore, we empirically introduce a threshold $\tau$ where we only minimize the message retrieval loss when $t\leq\tau$. This strategy helps effectively inject the watermark while stabilizing the fine-tuning procedure.
Preserving Image Consistency
In addition to fingerprinting the output, it is critical to preserve the generated images’ quality after watermarking. In other words, the generated images after watermarking across different users should be visually equivalent. To achieve this, we treat the original model as an oracle and align the generated output with the original output. Since the whole generation involves multiple denoising steps, we need to ensure the $\epsilon_{\theta}(\hat{\mathbf{x}}_{t};w_{i})$ and $\epsilon_{\theta}(\mathbf{x}_{t})$ are aligned for each $t$. Therefore, we introduce the second optimization objective by minimizing over the image consistency loss, denoted as follows:
$\min_{\theta}\mathcal{L}_{c}(\epsilon_{\theta}(\hat{\mathbf{x}}_{t};w_{i}),\epsilon_{\theta_{ori}}(\mathbf{x}_{t})),$ (6)
where $\theta_{ori}$ represents the original diffusion model weights before watermarking and $\mathcal{L}_{c}$ is the Mean-Squared Loss. To further improve the image consistency between different watermarks, for $t>\tau$, we replace the conditioned input $w_{i}$ with a never-used null watermark denoted as $w_{null}$. By fixing the condition with $w_{null}$, we ensure that $\epsilon_{\theta}(\hat{\mathbf{x}}_{t};w_{null})$ is independent of $w_{i}$ and keeps unchanged until $t\leq\tau$, which yields an improved image consistency among diverse users. We demonstrate watermarked samples of our method along with Tree-Ring [32] in Figure 2 and can observe that our method significantly improves the image consistency between watermarked contents. From the results, our watermarked outputs are not only visually equivalent across different users but also maintain the same semantic meaning with only a slight visual difference.

### v2

@@ Abstract
The ethical need to protect AI-generated content has been a significant concern in recent years. While existing watermarking strategies have demonstrated success in detecting synthetic content (detection), there has been limited exploration in identifying the users responsible for generating these outputs from a single model (owner identification). In this paper, we focus on both practical scenarios and propose a unified watermarking framework for content copyright protection within the context of diffusion models. Specifically, we consider two parties: the model provider, who grants public access to a diffusion model via an API, and the users, who can solely query the model API and generate images in a black-box manner. Our task is to embed hidden information into the generated contents, which facilitates further detection and owner identification. To tackle this challenge, we propose a Watermark-conditioned Diffusion model called WaDiff, which manipulates the watermark as a conditioned input and incorporates fingerprinting into the generation process. All the generative outputs from our WaDiff carry user-specific information, which can be recovered by an image extractor and further facilitate forensic identification. Extensive experiments are conducted on two popular diffusion models, and we demonstrate that our method is effective and robust in both the detection and owner identification tasks. Meanwhile, our watermarking framework only exerts a negligible impact on the original generation and is more stealthy and efficient in comparison to existing watermarking strategies. Our code is publically available at https://github.com/rmin2000/WaDiff.
@@ 1 Introduction
Figure 1: Illustration of our proposed WaDiff. All users access the diffusion model by querying the public API and are assigned a unique watermark. The generation process is conditioned on the watermark, and each user’s generated outputs would carry specific fingerprinting information which is further utilized to identify the owner of the generated image.
@@
To enable the traceability of diffusion-generated images, a commonly employed strategy is to embed a unique fingerprint to contents generated by an individual user and then forensically identify the owner from the watermarked image. A line of previous works [1, 26, 12, 29, 45] has extensively investigated this approach within the realm of traditional multimedia copyright protection, which is commonly referred to as post-hoc watermark. Typically, these strategies involve embedding an imperceptible fingerprint into the generative content while leaving an identifiable trace that can be detected using a pre-designed mechanism. However, the post-hoc watermark requires additional computational costs for watermark injection and is more susceptible to circumvention. For instance, in the event of model leakage, attackers can easily detect and bypass the post-processing module. Recently, the Stable Signature [11] investigated the latent diffusion paradigm and proposed a method that incorporates watermarks into each latent decoder. However, solely fingerprinting the latent decoder limits their application scenario to the latent diffusion paradigm only and can be easily circumvented by retraining the latent decoder on a clean dataset. At the same time, since only a fixed watermark could be embedded into the latent decoder, every model has to be fine-tuned before being distributed to the users, making it hard to use in large-scale real-world systems. Considering both the scalability and effectiveness, we aim to investigate whether we can embed fingerprints during the generation process to incorporate user-specific watermarks without customized fine-tuning.
@@
In this paper, we introduce WaDiff (as shown in Figure 1), a watermark-conditioned diffusion model, which incorporates the watermark as a conditioned input and generates images with unique fingerprints tailored to individual users. Unlike previous approaches [11, 36], which solely focus on fingerprinting the latent decoder in the diffusion model, our watermarking strategy is seamlessly integrated into the diffusion generation process. This makes our approach more general and applicable to other types of diffusion models [14] that do not include a latent decoder, while simultaneously eliminating the need for post-processing. Our WaDiff builds upon a pre-trained diffusion model, with slight modifications to the original input layers to accommodate the inclusion of watermark information by expanding the channels. Specifically, we first embed the watermark bits through a linear layer and then concatenate this projected vector with the original input to construct the watermark-conditioned input. We design a unified watermarking framework with two novel objective functions, named message retrieval loss and consistency loss. The message retrieval loss ensures the effective embedding of fingerprints into the generated content, allowing for successful retrieval of the embedded watermark. Meanwhile, the consistency loss ensures that the inclusion of the watermark has a negligible effect on the overall generation quality.
@@
To this end, we evaluate our method on two popular open-source diffusion models and perform both detection and identification tasks. We also make detailed comparisons with widely used post-hoc watermarking strategies and Tree-Ring [34], a training-free fingerprinting framework for diffusion models. Experimental results demonstrate that our efficient watermarking strategy enables accurate detection and identification in a large-scale system with numerous users and remains robust across various data augmentations. Moreover, the generated images after watermarking maintain exceptional generation quality and are visually indistinguishable across different watermarks. Our contributions are organized as follows:
@@
  • Different from previous work, we propose a scalable watermarking strategy that efficiently integrates user-specific fingerprints into the diffusion generation process without the need for customized fine-tuning.
@@
  • We introduce WaDiff, a watermark-conditioned diffusion model, along with a unified watermarking framework. Unlike post-hoc fingerprinting, WaDiff manipulates the watermark as conditioned input and allows for effective watermarking while having a negligible impact on the generation quality.
@@
  • Extensive experiments demonstrate that our watermarking framework achieves precise and robust performance in detecting AI-generated content and identifying the source owner of generated images.
@@ 2.1 Diffusion Models
Diffusion models have recently shown tremendous capability for high-quality image generation [14, 30, 28, 27, 31]. Depending on the space in which the generation process is performed, diffusion models can be divided into pixel-space [14, 9] and latent-space diffusion models [27]. In pixel-space diffusion models, images are directly generated starting from sampled Gaussian noise [14]. To reduce the computational complexity, latent diffusion models have been proposed to first generate latent features from noise and then decode the latent features to images by VAE [27]. With the powerful generation capability, they have demonstrated amazing results on various computer vision tasks, such as text-to-image generation [23, 28], sketch-to-image generation [22], text-guided image editing [23].
@@ 2.2 Image Watermarking
Image watermarking has been a broadly explored technique for decades. Traditional strategies typically start from a host image and inject watermark information directly in the spatial domain [26, 29], or through certain domain transformations such as Discrete Cosine Transform (DCT) [1] and Discrete Wavelet Transform (DWT) [12]. With advancements in deep learning, researchers have also explored the use of deep-learning techniques to replace manually designed watermark patterns. For example, a line of works [45, 19, 32, 44, 40] leverages the capability of Deep Neural Networks (DNNs) to improve the stealthiness and robustness of the watermark. In addition to facilitating image protection, image watermarking techniques have been proved useful in various security challenges, such as model ownership verification [42, 35, 41], backdoor attack [17], dataset copyright protection [16], and forensic adversarial defense [4]. In our method, we leverage image watermarking to embed unique watermark information for individual users, thereby facilitating subsequent detection and identification.
@@ 2.3 Fingerprinting in Diffusion Models
In light of recent advancements in generative AI, researchers have been exploring watermarks to safeguard or regulate the usage of generative models [39, 38, 15, 43, 7, 21, 10, 5, 37]. In this study, we primarily focus on fingerprinting the generative content. The Stable Signature [11] first proposed a watermarking scheme for latent diffusion models [27] by fingerprinting a set of latent decoders. They then distributed these customized decoders to individual users for both detection and identification. Building upon this work, [36] improved the scalability by incorporating a message matrix into the latent decoder, allowing multiple watermarks to be carried within a single architecture. However, their methods only focus on manipulating the latent decoder while keeping the diffusion process intact, which makes their methods only applicable to latent-space diffusion models. Also, this leaves them vulnerable to attacks like simply retraining the fingerprinted latent decoder on a clean dataset. In a parallel work called Tree-Ring [34], researchers concealed the fingerprint within the frequency domain of the initial noisy vector. Detection is then performed by reversing the generative image back to the noisy vector and comparing it with the original noisy pattern. Despite its training-free nature, the detection process heavily relies on the time-consuming reversion process and poses challenges for identification among multiple users.
In contrast to previous studies, we propose the WaDiff along with a simple and unified watermarking framework for both detection and owner identification. Our method seamlessly integrates the fingerprinting process into the image generation process, which can be applied with various diffusion types and sampling schedules, resulting in enhanced stealthiness and compatibility.
@@ 3 Problem Setting
The model provider deploys a generation model and grants public access to $m$ registered users. Considering the intellectual property (IP) protection, the model architecture and parameters are concealed from users, which means only black-box access for users is permitted and all the internal information remains encrypted. Note that the black-box scenario is common in practice, and has been widely adopted in current generative models such as the Midjourney and ChatGPT [3]. Each user denoted as $u_{i}$, is required to register before usage and would be then assigned a unique ID $i$ where $i\in\{1,\dots,m\}$, representing the user’s unique identity. One of the users $u_{i}$ would use the provided generation model to generate a picture $\mathbf{p}$. In practice, the $\mathbf{p}$ might be further processed under several image manipulations $f$ such as resizing and compression. We aim to build a security auditing system to enable the detection and tracing of the diffusion-generated content $\mathbf{p}$.
We divided our task into two challenges. The first challenge involves determining whether the generative content originates from our diffusion model, which we refer to as detection. However, simple detection of the generative content is insufficient to differentiate who generated a particular image when multiple users are engaged. Therefore, our second challenge focuses on a more complex scenario, which entails accurate identification of the owner of $f(\mathbf{p})$, from a pool of users. To address both challenges effectively, we propose a unified watermarking framework named WaDiff, which incorporates the embedding of user-specific fingerprints within the generation process. In other words, the generative images of each user are watermarked with unique information that distinguishes them from other users and facilitates both detection and identification.
@@ 4.2 Rooting Fingerprints in Diffusion Process
Pre-training Watermark Decoders.
@@
Inspired by [11], we first train an image encoder $\mathcal{W}$ and decoder $\mathcal{D}$ such that $\mathcal{D}$ could retrieve the message $\mathbf{w}$ pre-embedded by $\mathcal{W}$. The $\mathcal{W}$ is then discarded and only the pre-trained $\mathcal{D}$ is left to serve as a reference to fine-tune the diffusion model. Specifically, the encoder $\mathcal{W}$ takes image $\mathbf{x}$ and $n$-bit binary message $\mathbf{w}\in\{0,1\}^{n}$ as input and encodes $\mathbf{w}$ as an imperceptible residual $\delta$ added to $\mathbf{x}$, where the extractor $\mathcal{D}$ aims to restore the pre-encoded $\mathbf{w}$ from the watermarked input $\mathcal{W}(\mathbf{x})=\mathbf{x}+\delta$. To train $\mathcal{W}$ and $\mathcal{D}$ in an end-to-end manner, we utilize two loss terms to fulfill the optimization objective. First, to accurately restore $\mathbf{w}$ from $\mathcal{W}(\mathbf{x})$, we minimize the Binary Cross Entropy (BCE) between the decoded output $\mathcal{D}(\mathcal{W}(\mathbf{x}))$ and ground-truth $\mathbf{w}$. Second, to reduce the visibility of $\delta$, we penalize $\|\delta\|_{2}$ such that the added message perturbation is less perceptible. We further notice that the $\mathcal{D}$ obtained with the above training schedule is sensitive to several data manipulations such as resizing and compression. Therefore, similar to the simulation layer in [11], we transform $\mathcal{W}(\mathbf{x})$ with random data augmentations during the training stage to improve the robustness of the watermark system. We formulate the complete training objective as follows:
$\min_{\mathcal{W},\mathcal{D}}\mathbb{E}_{\mathbf{x},\mathbf{w},f}[\mathcal{L}_{BCE}(\mathcal{D}(f(\mathcal{W}(\mathbf{x}))),\mathbf{w})+\gamma\|\delta\|_{2}],$ (3)
@@
where $\gamma>0$ is a hyperparameter to control the visibility of $\delta$ and $f$ is a randomly selected image transformation from a pool of data augmentations.
Watermark-conditioned Diffusion Model.
@@
In contrast to post-hoc strategies, where the fingerprinting process occurs after the entire generation process, our approach imprints fingerprints during the sampling process. To discern the source user of generated outputs, we assign each user $u_{i}$ a unique $\mathbf{w}_{i}$ as the conditioned input of the diffusion model such that $\mathbf{w}_{i}$ could be embedded to the generated content during inference and correctly restored by the pre-trained $\mathcal{D}$.
@@
To incorporate $\mathbf{w}_{i}$ into the generation process, we expand the input channels and conceal the watermarking information at each denoising step $t$. This is achieved by applying a linear layer to project $\mathbf{w}_{i}$ into $\mathcal{P}(\mathbf{w}_{i})\in\mathbb{R}^{\widetilde{C}\times H\times W}$, where $\widetilde{C}$ denotes the number of watermark channels in the latent space. We then concatenate it with the original latent variable $\mathbf{x}_{t}\in\mathbb{R}^{C\times H\times W}$ along their first dimension, resulting in the conditioned latent variable $\mathbf{\hat{x}}_{t,i}=concat(\mathbf{x}_{t},\mathcal{P}(\mathbf{w}_{i}))\in\mathbb{R}^{(C+\widetilde{C})\times H\times W}$. Note that channel expansion is a common manipulation to integrate additional information and has been widely used in previous works [45, 32]. The next step involves embedding $\mathbf{w}_{i}$ into the generative content to allow for detection by the pre-trained decoder $\mathcal{D}$. However, as the sampling process typically involves multiple denoising steps to obtain the generative output, it is challenging to directly incorporate it into the fine-tuning process. To address this issue, we start by restoring the original image $\mathbf{x}_{0}$ within a single step. Revisiting the forward noising in Equation 2, $\mathbf{x}_{0}$ could be directly recovered from $\mathbf{x}_{t}$ by subtracting the second noise term and subsequently rescaling. In the absence of the ground truth $\epsilon$, we take the prediction from the diffusion model as an estimate. In our watermark-conditioned model, we replace $\epsilon$ with $\epsilon_{\theta}(\hat{\mathbf{x}}_{t,i})$ and construct the conditioned reverse of $\mathbf{x}_{0}$ at time step $t$ as:
$\mathbf{x}_{0,i}^{t}=\frac{\mathbf{x}_{t}-\sqrt{1-\overline{\alpha}_{t}}\epsilon_{\theta}(\hat{\mathbf{x}}_{t,i})}{\sqrt{\overline{\alpha}_{t}}}.$ (4)
@@
$\min_{\theta}\mathbb{E}_{\mathbf{x},i,t}[\mathcal{L}_{m}(\mathcal{D}(\mathbf{x}_{0,i}^{t}),\mathbf{w}_{i})].$ (5)
By optimizing Equation 5, we ensure that the pre-embedded $\mathbf{w}_{i}$ in $\mathbf{x}_{0,i}^{t}$ could be detected by $\mathcal{D}$. However, the image quality of $\mathbf{x}_{0,i}^{t}$ is significantly influenced by the denoising step $t$, resulting in higher image quality for smaller $t$ and noisier images for larger $t$. Therefore, we empirically introduce a threshold $\tau$ where we only minimize the message retrieval loss when $t\leq\tau$. This strategy helps effectively inject the watermark while stabilizing the fine-tuning procedure.
Preserving Image Consistency.
In addition to fingerprinting the output, it is critical to preserve the generated images’ quality after watermarking. In other words, the generated images after watermarking across different users should be visually equivalent. To achieve this, we treat the original model as an oracle and align the generated output with the original output. Since the whole generation involves multiple denoising steps, we need to ensure the $\epsilon_{\theta}(\hat{\mathbf{x}}_{t,i})$ and $\epsilon_{\theta_{ori}}(\mathbf{x}_{t})$ are aligned for each $t$, where $\theta_{ori}$ represents the pre-trained diffusion model weights. Therefore, we introduce the second optimization objective by minimizing over the image consistency loss, denoted as follows:
$\min_{\theta}\mathbb{E}_{\mathbf{x},i,t}[\mathcal{L}_{c}(\epsilon_{\theta}(\hat{\mathbf{x}}_{t,i}),\epsilon_{\theta_{ori}}(\mathbf{x}_{t}))],$ (6)
where $\mathcal{L}_{c}$ is the Mean-Squared Loss. To further improve the image consistency between different watermarks, for $t>\tau$, we replace the conditioned input $\mathbf{w}_{i}$ with a never-used null watermark denoted as $\mathbf{w}_{null}$. By fixing the condition with $\mathbf{w}_{null}$, we ensure that $\epsilon_{\theta}(\hat{\mathbf{x}}_{t,null})$ is distinct from $\mathbf{w}_{i}$ and keeps unchanged until $t\leq\tau$, which yields an improved image consistency among diverse users. We demonstrate watermarked samples of our method along with Tree-Ring [34] in Figure 2 and can observe that our method significantly improves the image consistency between watermarked contents. From the results, our watermarked outputs are not only visually equivalent across different users but also maintain the original semantic meaning with only a slight visual difference.